OpenStudy (anonymous):

Which is equivalent to i^224?

OpenStudy (freckles):

what is the remainder when dividing 224 by 4?

OpenStudy (freckles):

\[i^{224}=i^{4(q)+r}=(i^4)^qi^r=(1)^qi^r=1i^r=i^r\]

OpenStudy (freckles):

where r is the remainder

OpenStudy (mathmath333):

224=4*56

OpenStudy (anonymous):

so it's 1?

OpenStudy (mathmath333):

\(\Huge\tt \begin{align} \color{black}{\checkmark}\end{align}\)

OpenStudy (solomonzelman):

yes, it is 1. 224 is divisible by 4:)